OpenStudy (anonymous):

What are the coordinates of Q, if the midpoint of is (2, 4) and the coordinates of P are (5, 9)?

OpenStudy (lgbasallote):

\[\huge \text{midpoint} \; = \; (\frac{x_1 + x_2}{2} \; , \; \frac{y_1 + y_2}{2})\] \[\huge (2, 4) = (\frac{5 + x}{2} \; , \; \frac{9 + y}{2})\] solve for x first \[2 = \frac{5+x}{2}\] then solve for y \[4 = \frac{9+ y}{2}\] do you get it?

OpenStudy (anonymous):

some what

OpenStudy (lgbasallote):

nice. so tell me what you get for x and y